OpenStudy (anonymous):

what is the restriction on the quotient of 5x^2-10x-15/x^2=9 divided buy 6x+12/x+2

OpenStudy (anonymous):

i think you wrote x^2-9

OpenStudy (anonymous):

i meant x+2 in the end part

OpenStudy (anonymous):

x cant be -4

OpenStudy (anonymous):

please repeat your statement once more.

OpenStudy (anonymous):

5x^2-10x-15/x^2-9 divided buy 6x+12/x+2

OpenStudy (anonymous):

\[\frac{ 5x ^{2}-10x-15 }{ x ^{2}-9 }\times \frac{ x+2 }{ 6x+12 }\] now you can solve. Find the values for denominator is 0.

OpenStudy (anonymous):

-3? @surjithayer

OpenStudy (anonymous):

x^2-9=(x+3)(x-3)

OpenStudy (anonymous):

does that mean its positive instead?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

3,-3,-2